# Towards Multimodal Human Intention Understanding Debiasing via Subject-Deconfounding

### v1

@@ Abstract
Multimodal intention understanding (MIU) is an indispensable component of human expression analysis (e.g., sentiment or humor) from heterogeneous modalities, including visual postures, linguistic contents, and acoustic behaviors. Existing works invariably focus on designing sophisticated structures or fusion strategies to achieve impressive improvements. Unfortunately, they all suffer from the subject variation problem due to data distribution discrepancies among subjects. Concretely, MIU models are easily misled by distinct subjects with different expression customs and characteristics in the training data to learn subject-specific spurious correlations, significantly limiting performance and generalizability across uninitiated subjects. Motivated by this observation, we introduce a recapitulative causal graph to formulate the MIU procedure and analyze the confounding effect of subjects. Then, we propose SuCI, a simple yet effective causal intervention module to disentangle the impact of subjects acting as unobserved confounders and achieve model training via true causal effects. As a plug-and-play component, SuCI can be widely applied to most methods that seek unbiased predictions. Comprehensive experiments on several MIU benchmarks clearly demonstrate the effectiveness of the proposed module.
Keywords:
Human intention understanding Multimodal learning
1 Introduction
“You must never let the appearance of any subject affect your judgment.”
Sherlock Holmes, The Sign of the Four
Figure 1: Examples on the MOSI benchmark [61] illustrate the subject variation problem. Multimodal expressions from four subjects potentially convey distinct semantic correlations due to their different customs and styles in expressing sentiments. Training with samples collected from the subjects (1, 2, 3) invariably results in the SOTA model [21] learning spurious connections between subject-specific semantics and negative sentiment polarity. In this case, the differences between the subject-specific relations captured via the multimodal manner and the characteristics of subject 4 would lead to the subject-related prediction bias.
@@ 1 Introduction
As a research hotspot that combines linguistic [3] and non-verbal behaviors (e.g., visual [6] and acoustic [37] modalities), multimodal intention understanding (MIU) has attracted significant attention from computer vision [4, 10, 48], natural language processing [35, 41], and speech recognition communities [49, 52] in recent years. Thanks to the progressive development of multimodal language benchmarks [13, 60, 61], extensive studies [36, 12, 27, 59, 58, 42, 43, 23, 28, 9, 34, 38, 18] have presented impressive multimodal models on training data containing distinct subjects, diverse topics, and different modalities. Despite the achievements of previous approaches by exploiting representation learning architectures [48, 9, 23] and fusion strategies [59, 42, 28], they invariably suffer from a prediction bias when applied to testing samples of new subjects.
@@
The harmful prediction bias is mainly caused by the subject variation problem. Specifically, different subjects’ expression styles and behaviors (e.g., facial expressions or acoustic information) from the training data are highly idiosyncratic in social communication, affected by the subjects’ customs or culture [20]. Once well-designed models are trained on such data, subject-specific semantic correlations (e.g., particular facial action unit co-occurrence [5]) would inexorably affect performance and generalizability. Worse still, the spurious connections between the trained models and specific subjects will be transmitted via multiple modalities when the data paradigm is extended from isolated to multimodal situations. Recall the prominent MIU benchmarks (e.g., MOSI [61] for sentiment analysis or UR_FUNNY [13] for humor detection), whose collectors advocated video-level data splitting so that segments from the same video will not appear across train, valid, and test splits. Although the trained models may avoid memorizing the average affective state of a subject [22], they cannot generalize well across uninitiated subjects. The examples in Fig. 1 provide strong evidence of this. Concretely, subjects 1, 2, and 3 tend to use sentimentally unimportant words “but” and “just” to express negative emotions. In this case, the state-of-the-art (SOTA) model [21] is misled to focus on spurious clues from the textual utterances of the subjects and make an entirely incorrect prediction when applied to subject 4. Similar observations are found in the visual and audio modalities. For instance, the trained model erroneously takes subject-specific facial appearances (e.g., “grimace and pursed mouth” from subjects 1 and 2) and acoustic behaviors (e.g., “agitated tone” from subjects 2 and 3) as semantic shortcuts to infer frustrating negative sentiment polarity.
@@
Motivated by the above observations, this paper aims to improve MIU methods by causal demystification rather than beating them. We propose a Subject Causal Intervention module (SuCI) to disentangle the impact of semantic differences among subjects from multimodal expressions. Specifically, we first formulate a universal causal graph to analyze the MIU procedure and interpret the causalities among the variables. From the causal inference perspective, the subjects are essentially regarded as confounders [11], which mislead the models to learn subject-specific semantic correlations in the training data, as well as causing prediction bias for new subjects in the inference phase. For clarity, we represent the multimodal inputs as $\bm{X}$ (cause) and its corresponding predictions as $\bm{Y}$ (effect). Existing MIU models aim to approximate $P(\bm{Y}|\bm{X})$ as much as possible while failing to perform unbiased predictions. Unlike conventional likelihood estimation $P(\bm{Y}|\bm{X})$, our SuCI achieves subject de-confounded training and removes subject-caused confounding effects by embracing causal intervention $P(\bm{Y}|do(\bm{X}))$ regarding the backdoor adjustment [30]. As a model-agnostic component, SuCI can be readily integrated into most MIU models to perform unbiased predictions by pursuing true causal effects among variables. We comprehensively evaluate the effectiveness of SuCI on several MIU benchmarks. Numerous experiments show that SuCI can significantly and consistently improve existing baselines, achieving new SOTA performance.
@@
To sum up, this paper has the following contributions. (i) We are the first to investigate the subject variation problem in MIU tasks via a tailored causal graph and identify the subjects as confounders which misleads the models to capture subject-specific spurious correlations and cause prediction bias. (ii) Based on the causal theory of backdoor adjustment, we present SuCI, a subject causal intervention module to remove prediction bias and confounding effects of subjects. (iii) The causality-based SuCI is general and suitable for various models with distinct structures and fusion strategies.
2 Related Work
2.1 Multimodal Intention Understanding
@@
Benefiting from available human communication resources and data, MIU benchmarks [60, 61, 13, 4] with different scales and typologies have been increasingly developed and applied in recent years. Recent MIU tasks focus on subject-centered intention understanding and behavior analysis from text, visual, and audio modalities, including but not limited to emotion recognition [28], sentiment analysis [12], and humor detection [13]. Considering the heterogeneous and asynchronous nature of multimodal languages, numerous works have presented seminal network structures [34, 23, 58, 38], fusion strategies [42, 43, 59, 36], and representation learning paradigms [48, 52, 9, 50, 19]. For instance, Zadeh et al[59] proposed a tensor fusion to capture uni-modal, bi-modal, and tri-modal interactions explicitly. Tsai et al[42] designed a multimodal transformer structure to learn an adaptive process from one modality to another to capture cross-modal element dependencies. Afterward, Yang et al[52] presented a factorized representation strategy to learn similarities and differences among multimodal languages. Despite substantial achievements, existing methods invariably suffer from performance bottlenecks due to subject-related prediction bias caused by the subject variation problem. In this paper, we identify subjects in MIU benchmarks as harmful confounders from a causal perspective and improve different models with the proposed SuCI.
2.2 Causal Demystification
@@
Causal demystification as a revolutionary statistical theory aims to pursue causal effects among observed variables rather than their shallow correlations. Benefiting from the advances in learning-based technologies [55, 51, 54, 53, 25, 26], several studies exploring causality are mainly divided into two channels: causal intervention [24, 57] and counterfactual reasoning [35, 40]. Intervention [30] focuses on altering the natural tendency of the independent variable to vary with other variables to eliminate the impact of adverse effects. Counterfactuals depict imagined outcomes produced by factual variables under different treatment conditions [31]. Given the unbiased estimations provided by causal inference, it is increasingly applied to various downstream tasks such as computer vision [56, 47], natural language processing [62, 15], and reinforcement learning [7]. Recently, Sun et al[39] revealed the harmful associations of text semantics in multimodal sentiment analysis through counterfactual manipulation. In contrast, we address the confounding effect from multiple modalities by embracing causal intervention, which is more generalizable and adaptable for multimodal approaches.
3 Methodology
3.1 Structural Causal Graph in MIU Tasks
@@
Figure 2: The proposed causal graph explains the causal effects of the MIU procedure. Nodes denote variables and arrows denote the direct causal effects. (a) The conventional likelihood estimation $P(\bm{Y}|\bm{X})$. (b) The causal intervention $P(\bm{Y}|do(\bm{X}))$.
@@
To systematically diagnose the confounding effect present in MIU tasks, we first design a recapitulative causal graph to summarize the causal relationships among variables. The mainstream graphical notation following the structured causal model [32] is adopted due to its intuitiveness and interpretability. Concretely, a causal graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$ is considered a directed acyclic graph, which represents how a set of variables $\mathcal{N}$ convey causal effects through the causal links $\mathcal{E}$. As shown in Fig. 2(a), there are four variables in the MIU causal graph, including multimodal inputs $\bm{X}$, multimodal representations $\bm{M}$, unintended confounders $\bm{Z}$, and predictions $\bm{Y}$. Note that our causal graph applies to various MIU approaches since it is highly general, imposing no constraints on the detailed implementations. The detailed causalities are described as follows.
@@
According to the causal theory [30], the confounders $\bm{Z}$ are the common cause of $\bm{X}$ and corresponding predictions $\bm{Y}$. The positive effect of the subject-agnostic multimodal semantics provided by $\bm{M}$ follows the desired causal path $\bm{X}\rightarrow\bm{M}\rightarrow\bm{Y}$, which we aim to achieve and pursue. Unfortunately, $\bm{Z}$ causes subject-related prediction bias and misleads trained models to learn subject-specific misleading semantics rather than pure causal effects, leading to biased predictions on uninitiated subjects. The detrimental effects follow the backdoor paths $\bm{X}\leftarrow\bm{Z}\rightarrow\bm{Y}$ and $\bm{X}\leftarrow\bm{Z}\rightarrow\bm{M}\rightarrow\bm{Y}$.
3.2 Causal Intervention via Backdoor Adjustment
Following the causal graph in Fig. 2(a), the MIU model relies on the likelihood $P(\bm{Y}|\bm{X})$ for predictions that suffer from backdoor effects. $P(\bm{Y}|\bm{X})$ can be decomposed by the Bayes rule as follows:
@@ 3.2 Causal Intervention via Backdoor Adjustment
where $\mathcal{F}_{m}(\cdot)$ denotes any vanilla MIU model to learn the multimodal representations $\bm{M}$. $\bm{z}$ is a stratum of confounders (i.e., a subject), which introduces the observational bias via $P(\bm{z}|\bm{X})$. Theoretically, an ideal solution would be to collect massive data samples to ensure that subjects with different expression characteristics are included in the training and testing sets. However, this way is unrealistic due to human-bounded rationality and social ethics issues [17]. To address this, we embrace causal intervention $P(\bm{Y}|do(\bm{X}))$ to interrupt the adverse effects propagating between $\bm{X}$ and $\bm{Y}$ along the backdoor paths via the backdoor adjustment theory [30]. The $do(\cdot)$ operator is an efficient approximation to implement the empirical intervention [11]. In our case, backdoor adjustment means measuring the causal effect of each stratum in the subject confounders and then performing a weighted integration based on the prior proportions of samples from different subjects in the training data to estimate the average causal effect. From  Fig. 2(b), the impact from $\bm{Z}$ to $\bm{X}$ is cut off since the model would enable the subject prototype as the confounder in each stratum to contribute equally to the predictions $\bm{Y}$ by $P(\bm{Y}|do(\bm{X}))$. Eq. 1 with the intervention is formulated as follows:
@@
Intuitively, the model is no longer disrupted by subject-specific spurious correlations in backdoor paths since $\bm{z}$ no longer affects $\bm{X}$. $P(\bm{z})$ is the prior probability that depicts the proportion of each $\bm{z}$ in the whole.
Figure 3: A general MIU pipeline for the subject de-confounded training. The red dotted box shows the core component that achieves the approximation to causal intervention: our SuCI. SuCI can be readily integrated into the vanilla MIU model via backdoor adjustment to mitigate subject-specific spurious correlations and achieve debiased predictions in downstream MIU tasks.
3.3 Subject De-confounded Training with SuCI
We present a plug-in Subject Causal Intervention module (SuCI) to convert the theoretical intervention in Eq. 2 into a practical implementation. As Fig. 3 shows, SuCI can be readily integrated into the vanilla MIU model to estimate $P(\bm{Y}|do(\bm{X}))$ through the subject de-confounded training. The implementation details are as follows.
Subject-specific Feature Disentanglement. How to effectively disentangle subject-specific features from heterogeneous modalities is key to determining confounders. Considering that subject-related spurious semantics are usually distributed among different frames of multimodal sequences [48], we first devise a dynamic fusion mechanism to aggregate the frame-level semantic clues. Let the text, visual, and audio modality sequences from the corresponding feature extractors be denoted as $\bm{x}_{t}\in\mathbb{R}^{T_{t}\times d_{t}}$, $\bm{x}_{v}\in\mathbb{R}^{T_{v}\times d_{v}}$, and $\bm{x}_{a}\in\mathbb{R}^{T_{a}\times d_{a}}$, respectively, where $T_{(\cdot)}$ is the frame length and $d_{(\cdot)}$ is the embedding dimension. The dynamic fusion procedure is formulated as follows:
@@ 3.3 Subject De-confounded Training with SuCI
where [,][\,,][ , ] stands for the concatenation operator and $d_{s}=d_{t}+d_{v}+d_{a}$, The generator is implemented as a two-layer perceptron with the GeLU [14] activation. $\bm{s}$ is distilled by the following optimization objective:
@@
where $\mathcal{CE}(\cdot)$ and $\mathcal{MSE}(\cdot)$ represent the cross-entropy loss and mean squared error loss, respectively. $\mathcal{D}_{s}(\cdot;\theta_{\mathcal{D}_{s}})$ and $\mathcal{D}_{t}(\cdot;\theta_{\mathcal{D}_{t}})$ are the subject discriminator and task discriminator parameterized by $\theta_{\mathcal{D}_{s}}$ and $\theta_{\mathcal{D}_{t}}$, respectively, which consist of feed-forward neural layers. $y_{s}$ is the subject label determined by the dataset’s subject index. $C_{t}$ is the number of categories in downstream MIU tasks. In Eq. 6, $\bm{s}$ is supervised to encourage the output logits of the task discriminator to be equally distributed among all prediction categories to exclude task-related semantic information. Also, the subject discriminator is optimized to ensure that $\bm{s}$ belongs to a given subject, i.e., contains subject-specific semantics.
Confounder Construction. Confounder construction aims to make the model measure the causal effect of subject confounders among different strata during training to avoid subject-related prediction bias. Considering that subject features are similar within the same stratum and different across strata [30], we utilize all the features from a specific subject as the subject prototype, which represents the universal attributes of confounders in a specific stratum. Concretely, we maintain a memory cache for each subject during training to store and update the subject prototype as a confounder, which is computed as ${\bm{z}_{i}=\frac{1}{N_{i}}\textstyle\sum_{k=1}^{N_{i}}}\bm{s}_{k}^{i}$. $N_{i}$ is the number of training samples for the $i$-th subject, and $\bm{s}_{k}^{i}$ denotes the $k$-th feature of the $i$-th subject. $\bm{z}_{i}$ is updated at the end of each epoch. In practice, each $\bm{z}_{i}$ is initialized by the uniform distribution to ensure stable training at the first epoch. Based on the number of subjects $N_{c}$, a stratified confounder dictionary is constructed, which is formulated as $\bm{Z}=[\bm{z}_{1},\bm{z}_{2},\dots,\bm{z}_{N_{c}}]$.
Intervention Instantiation. Estimating $P(\bm{Y}|do(\bm{X}))$ in practice is high overhead since it requires the forward computation of $\bm{X}$ and $\bm{z}$ for each pair. To reduce the overhead, we introduce the Normalized Weighted Geometric Mean (NWGM) [46] to achieve feature-level approximation for intervention instantiation:
@@
In this case, causal intervention makes $\bm{X}$ contribute fairly to the predictions of $\bm{Y}$ by incorporating every $\bm{z}$. We parameterize a network model to approximate Eq. 7 as:
@@
where $\bm{W}_{m}\in\mathbb{R}^{d_{h}\times d}$ and $\bm{W}_{h}\in\mathbb{R}^{d_{h}\times d_{s}}$ are the learnable parameters. $\bm{m}\in\mathbb{R}^{d\times 1}$ is the multimodal representation produced by the vanilla MIU model. The above approximation is reasonable since the effect on $\bm{Y}$ is attributed to $\bm{M}$ and $\bm{Z}$ in the MIU causal graph. Then, $\mathbb{E}[h(\bm{z})]$ is approximated as an adaptive aggregation for all confounders according to the backdoor adjustment of Sec. 3.2, which is formulated as follows:
@@
4 Experiments
4.1 Benchmarks and Evaluation Metrics
Extensive experiments are conducted on three mainstream MLU benchmarks. Concretely, MOSI [61] is a multimodal human sentiment analysis dataset consisting of 2,199 video segments. The standard data partitioning is 1,284 samples for training, 284 samples for validation, and 686 samples for testing. These samples contain a total of 89 distinct subjects from video blogs. Each sample is manually annotated with a sentiment score ranging from -3 to 3, including highly negative, negative, weakly negative, neutral, weakly positive, positive, and highly positive. MOSEI [60] is a large-scale human sentiment and emotion recognition benchmark containing 22,856 video clips from 1,000 different subjects and 250 diverse topics. Among these samples, 16,326, 1,871, and 4,659 samples are used as training, validation and testing sets. The sample annotation protocols used for sentiment analysis are consistent with MOSI. UR_FUNNY [13] is a multimodal human humor detection dataset that contains 16,514 video clips from 1,741 subjects collected by the TED portal. There are 10,598, 2,626, and 3,290 samples in the training, validation, and testing sets. Each sample provides the target punchlines and associated contexts from multimodal utterances to support the detection of subject humor/non-humor in the binary labels.
Evaluation Metrics. Following widely adopted implementations [42, 48, 21], we use the 7-class accuracy ($Acc_{7}$), binary accuracy ($Acc_{2}$), and $F1$ score to evaluate the results on MOSI and MOSEI. The standard binary accuracy ($Acc_{2}$) is utilized for evaluation on UR_FUNNY.
4.2 Model Zoo
Considering applicability and scalability, we choose five reproducible models with entirely different network structures to verify the proposed plug-in SuCI. A brief overview is as follows. Specifically, MulT [42] constructs multimodal transformers to learn element dependencies between pairwise modalities. MISA [9] captures modality-invariant and modality-specific diverse representations based on feature disentanglement. Self-MM [58] utilizes a self-supervised paradigm to learn additional emotion semantics from unimodal label generation. MMIM [12] proposes a hierarchical mutual information maximization to alleviate the loss of valuable task-related clues. DMD [21] designs cross-modal knowledge distillations to bridge the semantic gap among heterogeneous modalities.

### v3

# Debiased Multimodal Understanding for Human Language Sequences

@@ Abstract
Human multimodal language understanding (MLU) is an indispensable component of expression analysis (e.g., sentiment or humor) from heterogeneous modalities, including visual postures, linguistic contents, and acoustic behaviours. Existing works invariably focus on designing sophisticated structures or fusion strategies to achieve impressive improvements. Unfortunately, they all suffer from the subject variation problem due to data distribution discrepancies among subjects. Concretely, MLU models are easily misled by distinct subjects with different expression customs and characteristics in the training data to learn subject-specific spurious correlations, limiting performance and generalizability across new subjects. Motivated by this observation, we introduce a recapitulative causal graph to formulate the MLU procedure and analyze the confounding effect of subjects. Then, we propose SuCI, a simple yet effective causal intervention module to disentangle the impact of subjects acting as unobserved confounders and achieve model training via true causal effects. As a plug-and-play component, SuCI can be widely applied to most methods that seek unbiased predictions. Comprehensive experiments on several MLU benchmarks clearly show the effectiveness of the proposed module.
Introduction
@@
As a research hotspot that combines linguistic and non-verbal behaviours (e.g., visual and acoustic modalities), human multimodal language understanding (MLU) has attracted significant attention from computer vision (Yang et al. 2022a), natural language processing (Tian et al. 2022), and speech recognition communities (Yang et al. 2022b, d) in recent years. Thanks to the progressive development of multimodal language benchmarks (Hasan et al. 2019; Zadeh and Pu 2018; Zadeh et al. 2016), extensive studies (Rahman et al. 2020; Han, Chen, and Poria 2021; Liu et al. 2018; Yu et al. 2021; Tsai et al. 2019; Liang et al. 2021b; Lv et al. 2021; Pham et al. 2019; Sun et al. 2022; Lei et al. 2023) have presented impressive multimodal models on training data containing distinct subjects, diverse topics, and different modalities. Despite the achievements of previous approaches by exploiting representation learning architectures (Yang et al. 2022a; Liang et al. 2021b) and fusion strategies (Tsai et al. 2019; Lv et al. 2021), they invariably suffer from a prediction bias when applied to testing samples of new subjects.
@@
The harmful prediction bias is mainly caused by the subject variation problem. Specifically, different subjects’ expression styles and behaviours (e.g., facial expressions or acoustic information) from the training data are highly idiosyncratic in social communication, affected by the subjects’ customs or culture (Li and Deng 2020). Once well-designed models are trained on such data, subject-specific semantic correlations (e.g., particular facial action unit co-occurrence (Chen et al. 2022)) would inexorably affect performance and generalizability. Worse still, the spurious connections between the trained models and specific subjects will be transmitted via multiple modalities when the data paradigm is extended from isolated to multimodal situations. Recall the prominent MLU benchmarks (e.g., MOSI (Zadeh et al. 2016) for sentiment analysis or UR_FUNNY (Hasan et al. 2019) for humor detection), whose collectors advocated video-level data splitting so that segments from the same video will not appear across train, valid, and test splits. Although the trained models may avoid memorizing the average affective state of a subject (Liang et al. 2021a), they cannot generalize well across new subjects. The examples in Figure 1 provide strong evidence of this. Concretely, subjects 1, 2, and 3 tend to use sentimentally unimportant words “but” and “just” to express negative emotions. In this case, the MLU model (Li, Wang, and Cui 2023) is misled to focus on spurious clues from the textual utterances of the subjects and make an entirely incorrect prediction when applied to subject 4. Similar observations are found in the visual and audio modalities. For instance, the trained model erroneously takes subject-specific facial appearances (e.g., “grimace and pursed mouth” from subjects 1 and 2) and acoustic behaviours (e.g., “agitated tone” from subjects 2 and 3) as semantic shortcuts to infer frustrating negative sentiment.
Figure 1: Examples on the MOSI benchmark illustrate the subject variation problem. Multimodal expressions from four subjects potentially convey distinct semantic correlations due to their different customs and styles in expressing sentiments.
@@
Motivated by the above observations, this paper aims to improve MLU methods by causal demystification rather than beating them. We propose a Subject Causal Intervention module (SuCI) to disentangle the impact of semantic differences among subjects from multimodal expressions. Specifically, we first formulate a universal causal graph to analyze the MLU procedure and interpret the causalities among the variables. From the causal inference perspective, the subjects are essentially regarded as confounders (Glymour, Pearl, and Jewell 2016), which mislead the models to learn subject-specific semantic correlations in the training data, as well as causing prediction bias for new subjects in the inference phase. For clarity, we represent the multimodal inputs as $\bm{X}$ (cause) and its corresponding predictions as $\bm{Y}$ (effect). Existing MLU models aim to approximate $P(\bm{Y}|\bm{X})$ as much as possible while failing to perform unbiased predictions. Unlike conventional likelihood estimation $P(\bm{Y}|\bm{X})$, our SuCI achieves subject de-confounded training and removes subject-caused confounding effects by embracing causal intervention $P(\bm{Y}|do(\bm{X}))$ regarding the backdoor adjustment rule (Pearl 2009a). As a model-agnostic component, SuCI can be readily integrated into most MLU models to perform unbiased predictions by pursuing true causal effects among variables.
@@
To sum up, this paper has the following contributions. (i) We investigate the subject variation problem in MLU tasks via a tailored causal graph and identify the subjects as confounders which misleads the models to capture subject-specific spurious correlations and cause prediction bias. (ii) Based on the causal theory of backdoor adjustment, we present SuCI, a subject causal intervention module to remove prediction bias and confounding effects of subjects. (iii) We evaluate the effectiveness of SuCI on several MLU benchmarks. Experimental results show that SuCI can significantly and consistently improve existing baselines, achieving new SOTA performance.
Related Work
@@
Human Multimodal Language Understanding. Benefiting from available human communication resources and data, MLU benchmarks (Zadeh and Pu 2018; Zadeh et al. 2016; Hasan et al. 2019) with different scales and typologies have been increasingly developed and applied in recent years. Recent MLU tasks focus on subject-centered intention understanding and behavior analysis from text, visual, and audio modalities, including but not limited to emotion recognition (Lv et al. 2021), sentiment analysis (Han, Chen, and Poria 2021), and humor detection (Hasan et al. 2019). Considering the heterogeneous nature of multimodal languages, numerous works have presented seminal network structures (Pham et al. 2019; Liang et al. 2021b; Yu et al. 2021; Sun et al. 2022), fusion strategies (Tsai et al. 2019, 2018; Zadeh et al. 2017; Rahman et al. 2020), and representation learning paradigms (Yang et al. 2022a, d, c; Li, Yang, and Zhang 2023). For instance, MFSA (Yang et al. 2022d) presented a factorized representation strategy to learn similarities and differences among multimodal languages. Despite achievements, existing methods invariably suffer from performance bottlenecks due to subject-related prediction bias. In comparison, we identify subjects in MLU benchmarks as harmful confounders from a causal perspective and improve different models with our SuCI.
@@
Causal Demystification. Causal demystification as a potential statistical theory aims to pursue causal effects among observed variables rather than their shallow correlations. Benefiting from the advances in learning-based technologies (Yang et al. 2023e, b, d, c; Liu et al. 2023a, b), several studies exploring causality are mainly divided into two channels: causal intervention (Lin et al. 2022; Yang et al. 2021) and counterfactual reasoning (Qian et al. 2021; Tang et al. 2020). Intervention (Pearl 2009a) focuses on altering the natural tendency of the independent variable to vary with other variables to eliminate the impact of adverse effects. Counterfactuals depict imagined outcomes produced by factual variables under different treatment conditions (Pearl 2009b). Given the unbiased estimations provided by causal inference, it is increasingly applied to various downstream tasks such as computer vision (Yang, Zhang, and Cai 2021; Yang et al. 2023a), natural language processing (Zhang et al. 2021; Huang, Liu, and Bowman 2020), and reinforcement learning (Dasgupta et al. 2019). In this paper, we address the confounding effect from multiple modalities by embracing causal intervention, which is more generalizable and adaptable for multimodal approaches.
Methodology
Structural Causal Graph in MLU Tasks
@@
Figure 2: The causal graph explains causal effects of MLU procedure. Nodes denote variables and arrows denote the direct causal effects. (a) The conventional likelihood estimation $P(\bm{Y}|\bm{X})$. (b) The causal intervention $P(\bm{Y}|do(\bm{X}))$.
@@
To systematically diagnose the confounding effect present in MLU tasks, we first design a recapitulative causal graph to summarize the causal relationships among variables. The mainstream graphical notation following the structured causal model (Pearl et al. 2000) is adopted due to its intuitiveness and interpretability. Concretely, a causal graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$ is considered a directed acyclic graph, which represents how a set of variables $\mathcal{N}$ convey causal effects through the causal links $\mathcal{E}$. As shown in Figure 2(a), there are four variables in the MLU causal graph, including multimodal inputs $\bm{X}$, multimodal representations $\bm{M}$, unintended confounders $\bm{Z}$, and predictions $\bm{Y}$. Note that our causal graph applies to various MLU approaches since it is highly general, imposing no constraints on the detailed implementations. The causalities are shown below.
@@
According to the causal theory (Pearl 2009a), the confounders $\bm{Z}$ are the common cause of $\bm{X}$ and corresponding predictions $\bm{Y}$. The positive effect of the subject-agnostic multimodal semantics provided by $\bm{M}$ follows the desired causal path $\bm{X}\rightarrow\bm{M}\rightarrow\bm{Y}$, which we aim to achieve and pursue. Unfortunately, $\bm{Z}$ causes subject-related prediction bias and misleads trained models to learn subject-specific misleading semantics rather than pure causal effects, leading to biased predictions on uninitiated subjects. The detrimental effects follow the backdoor paths $\bm{X}\leftarrow\bm{Z}\rightarrow\bm{Y}$ and $\bm{X}\leftarrow\bm{Z}\rightarrow\bm{M}\rightarrow\bm{Y}$.
Causal Intervention via Backdoor Adjustment
Following the causal graph in Figure 2(a), the MLU model relies on the likelihood $P(\bm{Y}|\bm{X})$ for predictions that suffer from backdoor effects, which can be decomposed by the Bayes rule as follows:
@@
where $\mathcal{F}_{m}(\cdot)$ denotes any vanilla MLU model to learn the multimodal representations $\bm{M}$. $\bm{z}$ is a stratum of confounders (i.e., a subject), which introduces the observational bias via $P(\bm{z}|\bm{X})$. Theoretically, an ideal solution would be to collect massive data samples to ensure that subjects with different expression characteristics are included in the training and testing sets. However, this way is unrealistic due to social ethics issues (Jones 1999). To address this, we embrace causal intervention $P(\bm{Y}|do(\bm{X}))$ to interrupt the adverse effects propagating between $\bm{X}$ and $\bm{Y}$ along the backdoor paths via the backdoor adjustment theory (Pearl 2009a). The $do(\cdot)$ operator is an efficient approximation to implement the empirical intervention (Glymour, Pearl, and Jewell 2016). In our case, backdoor adjustment means measuring the causal effect of each stratum in the subject confounders and then performing a weighted integration based on the prior proportions of samples from different subjects in the training data to estimate the average causal effect. From Figure 2(b), the impact from $\bm{Z}$ to $\bm{X}$ is cut off since the model would enable the subject prototype as the confounder in each stratum to contribute equally to the predictions $\bm{Y}$ by $P(\bm{Y}|do(\bm{X}))$. Eq. (1) with the intervention is formulated as:
@@
The model is no longer disrupted by subject-specific spurious correlations in backdoor paths since $\bm{z}$ no longer affects $\bm{X}$. $P(\bm{z})$ is the prior probability that depicts the proportion of each $\bm{z}$ in the whole.
Figure 3: A general MLU pipeline for the subject de-confounded training. The red dotted box shows the core component that achieves the approximation to causal intervention: our SuCI. SuCI can be readily integrated into the vanilla MLU model via backdoor adjustment to mitigate subject-specific spurious correlations and achieve debiased predictions in downstream tasks.
Subject De-confounded Training with SuCI
We present a plug-in Subject Causal Intervention module (SuCI) to convert the theoretical intervention in Eq. (2) into a practical implementation. As Figure 3 shows, SuCI can be readily integrated into the vanilla MLU model to estimate $P(\bm{Y}|do(\bm{X}))$ through the subject de-confounded training. The implementation details are as follows.
Subject-specific Feature Disentanglement. How to effectively disentangle subject-specific features from heterogeneous modalities is key to determining confounders. Considering that subject-related spurious semantics are usually distributed among different frames of multimodal sequences (Yang et al. 2022a), we first devise a dynamic fusion mechanism to aggregate the frame-level semantic clues. Let the text, visual, and audio modality sequences from the corresponding feature extractors be denoted as $\bm{x}_{t}\in\mathbb{R}^{T_{t}\times d_{t}}$, $\bm{x}_{v}\in\mathbb{R}^{T_{v}\times d_{v}}$, and $\bm{x}_{a}\in\mathbb{R}^{T_{a}\times d_{a}}$, respectively, where $T_{(\cdot)}$ is the frame length and $d_{(\cdot)}$ is the embedding dimension. The dynamic fusion procedure is formulated as follows:
@@
where [,][\,,][ , ] stands for the concatenation operator and $d_{s}=d_{t}+d_{v}+d_{a}$, The generator is implemented as a two-layer perceptron with GeLU (Hendrycks and Gimpel 2016) activation. $\bm{s}$ is distilled by the following objective:
@@
where $\mathcal{CE}(\cdot)$ and $\mathcal{MSE}(\cdot)$ represent the cross-entropy loss and mean squared error loss, respectively. $\mathcal{D}_{s}(\cdot;\theta_{\mathcal{D}_{s}})$ and $\mathcal{D}_{t}(\cdot;\theta_{\mathcal{D}_{t}})$ are the subject discriminator and task discriminator parameterized by $\theta_{\mathcal{D}_{s}}$ and $\theta_{\mathcal{D}_{t}}$, respectively, which consist of feed-forward neural layers. $y_{s}$ is the subject label determined by the dataset’s subject index. $C_{t}$ is the number of categories in downstream tasks. In Eq. (6), $\bm{s}$ is supervised to encourage the output logits of the task discriminator to be equally distributed among all prediction categories to exclude task-related semantic information. Also, the subject discriminator is optimized to ensure that $\bm{s}$ belongs to a given subject, i.e., contains subject-specific semantics.
Confounder Construction. Confounder construction aims to make the model measure the causal effect of subject confounders among different strata during training to avoid subject-related prediction bias. Considering that subject features are similar within the same stratum and different across strata (Pearl 2009a), we utilize all the features from a specific subject as the subject prototype, which represents the universal attributes of confounders in a specific stratum. Concretely, we maintain a memory cache for each subject during training to store and update the subject prototype as a confounder, which is computed as ${\bm{z}_{i}=\frac{1}{N_{i}}\textstyle\sum_{k=1}^{N_{i}}}\bm{s}_{k}^{i}$. $N_{i}$ is the number of training samples for the $i$-th subject, and $\bm{s}_{k}^{i}$ denotes the $k$-th feature of the $i$-th subject. $\bm{z}_{i}$ is updated at the end of each epoch. In practice, each $\bm{z}_{i}$ is initialized by the uniform distribution to ensure stable training at the first epoch. Based on the number of subjects $N_{c}$, a stratified confounder dictionary is constructed, which is formulated as $\bm{Z}=[\bm{z}_{1},\bm{z}_{2},\dots,\bm{z}_{N_{c}}]$.
Intervention Instantiation. Estimating $P(\bm{Y}|do(\bm{X}))$ in practice is high overhead since it requires the forward computation of $\bm{X}$ and $\bm{z}$ for each pair. To reduce the overhead, we introduce the Normalized Weighted Geometric Mean (NWGM) (Xu et al. 2015) to achieve feature-level approximation for intervention instantiation:
@@
In this case, causal intervention makes $\bm{X}$ contribute fairly to the predictions of $\bm{Y}$ by incorporating every $\bm{z}$. We parameterize a network model to approximate Eq. (7) as follows:
@@
where $\bm{W}_{m}\in\mathbb{R}^{d_{h}\times d}$ and $\bm{W}_{h}\in\mathbb{R}^{d_{h}\times d_{s}}$ are the learnable parameters. $\bm{m}\in\mathbb{R}^{d\times 1}$ is the multimodal representation produced by the vanilla MLU model. The above approximation is reasonable since the effect on $\bm{Y}$ is attributed to $\bm{M}$ and $\bm{Z}$ in the MLU causal graph. Then, $\mathbb{E}[h(\bm{z})]$ is approximated as an adaptive aggregation for all confounders according to backdoor adjustment, which is formulated as:
@@
Experiments